# A solution to the Cauchy dual subnormality problem for a cyclic analytic $2$-isometry with defect operator of rank two

### v1

@@ 2. Case (ii)
$a_{11}$ $=(a^{2}-b)(|1+\xi|^{2})+\displaystyle\frac{2b}{\Delta}$
$a_{12}$ $=-b(1+\xi)$
$a_{22}$ $=\displaystyle 1-3b+2ab-\frac{b}{\Delta}$
It can be seen that
$b=|\alpha_{1}\alpha_{2}|=5,~~~~a=\displaystyle\frac{3b}{b+1}=\frac{5}{2},~~~~|1+\xi|^{2}=\displaystyle\frac{16}{5}$
and
$\Delta=\left[(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}\right]=1.$
This gives,
$a_{11}=14,~~a_{12}=-8-4i~~~a_{22}=6.$
Though, we have $D(\mu)=H(B)$, we note here that we cannot apply the Theorem 3.2 since it requires $\alpha_{1}$ and $\alpha_{2}$ to be distinct. Therefore, we modify the condition for our case as follows:
We begin by stating a partial fraction formula which can be easily verified.
@@
We apply this to polynomials $p_{1}(z)/z$ and $p_{2}(z)/z$ to obtain:
@@
Note that
$\frac{z}{z-\alpha}=-\displaystyle\sum_{l=0}^{\infty}\frac{z^{l+1}}{\alpha^{l+1}}\text{ and }\frac{z}{(z-\alpha)^{2}}=\sum_{l=0}^{\infty}\frac{(l+1)z^{l+1}}{\alpha^{l+2}}$
@@
$b_{1}(z)$ $=\displaystyle\frac{p_{1}(z)}{q(z)}=\frac{p_{1}(z)}{(z-\alpha)^{2}}=\frac{p_{11}}{\alpha}\sum_{l=0}^{\infty}\frac{z^{l+1}}{\alpha^{l+1}}+\frac{p_{1}(\alpha)}{\alpha^{2}}\sum_{l=0}^{\infty}\frac{(l+1)z^{l+1}}{\alpha^{l+2}}$
$b_{2}(z)$ $=\displaystyle\frac{p_{2}(z)}{q(z)}=\frac{p_{2}(\alpha)}{\alpha^{2}}\sum_{l=0}^{\infty}\frac{(l+1)z^{l+1}}{\alpha^{l+2}}.$ (3.25)
Let $b_{j}(z)=\displaystyle\sum_{m=1}^{\infty}b_{j,m}z^{m}$, $z\in\mathbb{D}$ and $j=1,2$. We note here that $B_{m}=(b_{1,m},b_{2,m})$, $m\geq 1$.
Using equations (3.25) we get,
$b_{1,m}=\displaystyle\frac{1}{\alpha^{m+1}}\left(p_{11}+\frac{p_{1}(\alpha)m}{\alpha^{2}}\right),~~m\geq 1$
and
$b_{2,m}=\displaystyle\frac{p_{2}(\alpha)~m}{\alpha^{m+3}},~~m\geq 1.$
Thus, $\forall m\geq 0,s\geq 0$, we get,
$b_{1,m+1+s}=\displaystyle\frac{1}{\alpha^{m+2+s}}\left(p_{11}+\frac{p_{1}(\alpha)(m+1+s)}{\alpha^{2}}\right)$
and
$b_{2,m+1+s}=\displaystyle\frac{p_{2}(\alpha)~(m+1+s)}{\alpha^{m+4+s}}.$
Hence,
@@
$=\displaystyle\frac{1}{\alpha^{m+2+s}}\left(p_{11}+\frac{p_{1}(\alpha)}{\alpha^{2}}(m+1+s)\right)\frac{1}{\overline{\alpha}^{n+2+s}}\overline{\left(p_{11}+\frac{p_{1}(\alpha)}{\alpha^{2}}(n+1+s)\right)}$
$\hskip 14.22636pt+\displaystyle\frac{p_{2}(\alpha)~(m+1+s)}{\alpha^{m+4+s}}\times\overline{\frac{p_{2}(\alpha)~(n+1+s)}{\alpha^{n+4+s}}}$
$=\displaystyle\frac{1}{|\alpha|^{2s}}\frac{1}{\alpha^{m+2}\overline{\alpha}^{n+2}}\Big(|p_{11}|^{2}+\frac{|p_{1}(\alpha)|^{2}}{|\alpha|^{4}}(m+1+s)(n+1+s)$
$\hskip 14.22636pt+\displaystyle\frac{p_{11}\overline{p_{1}(\alpha)}}{\overline{\alpha}^{2}}(n+1+s)+\frac{\overline{p_{11}}p_{1}(\alpha)}{\alpha^{2}}(m+1+s)\Big)$
$\hskip 14.22636pt+\frac{|p_{2}(\alpha)|^{2}(m+1+s)(n+1+s)}{|\alpha|^{2s}\alpha^{m+4}\overline{\alpha}^{n+4}}$
@@
We will show that, for $l=1$, the $2\times 2$ principal minor of the matrix
$\sum_{s=0}^{l}(-1)^{s}\binom{l}{s}((B_{1,m+1+s}B^{*}_{n+1+s}))_{m,n\geq 0}$ (3.26)
is negative and thus by [16, Theorem 3.5] we can conclude that the Cauchy dual of $M_{z}$ on $H(B)$ is not subnormal.
@@
$N=\left|(n_{ij})_{0\leq i,j\leq 1}\right|$.
Hence,
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{4}\overline{\alpha}^{2}}\left(1-\frac{2}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{4}\alpha^{2}}\left(1-\frac{2}{|\alpha|^{2}}\right)$
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{4}\overline{\alpha}^{3}}\left(2-\frac{3}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{6}\alpha}\left(1-\frac{2}{|\alpha|^{2}}\right)$
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{6}\overline{\alpha}}\left(1-\frac{2}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{4}\alpha^{3}}\left(2-\frac{3}{|\alpha|^{2}}\right)$
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{6}\overline{\alpha}^{2}}\left(2-\frac{3}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{6}\alpha^{2}}\left(2-\frac{3}{|\alpha|^{2}}\right)$
@@
$p_{11}=\sqrt{14},~p_{22}=\displaystyle\frac{2}{\sqrt{14}},~~p_{1}(\alpha)=\frac{20-30i}{\sqrt{14}}~~p_{2}(\alpha)=\frac{6+8i}{\sqrt{14}}$
@@
$n_{00}=\displaystyle\frac{228}{625},~~~n_{01}=\frac{344+512i}{3125},~~~n_{10}=\frac{344-512i}{3125},~~~n_{11}=\frac{332}{3125}$
Thus, $N=\displaystyle\frac{-2000}{9765625}<0.$
Hence, by [16, Theorem 3.5] we conclude that the Cauchy dual of $M_{z}$ on $H(B)$ is not subnormal.
Therefore, whenever measure $\mu=\delta_{\zeta_{1}}+\delta_{\zeta_{2}}$ where $\zeta_{1}$ and $\zeta_{2}$ are two distinct non-antipodal points on $\mathbb{T}$, the Cauchy dual of the operator $M_{z}$ on corresponding $D(\mu)$ is not subnormal.
Finally, [16, Theorem 2.4] ensures that if $\mu$ is supported at antipodal points of $\mathbb{T}$, then the Cauchy dual of the operator $M_{z}$ on the corresponding $D(\mu)$ is subnormal. This completes the proof. ∎
@@ 4. Epilogue
The authors believe that the main theorem of the present article can be proved for the measure $\mu$ in the genral form, i.e. $\mu=c_{1}\delta_{\zeta_{1}}+c_{2}\delta_{\zeta_{2}},$ where $c_{1},c_{2}$ are positive real numbers, but the computational complexity increases substantially. In addition, it would be interesting to extend the main theorem when a measure $\mu$ supported on three or more points on the unit circle. In particular, a suitable generalization of the notion of points on the unit circle being antipodal needs to be explored.

### v2

@@ 2. Case (ii)
$a_{11}=(a^{2}-b)(|1+\xi|^{2})+\displaystyle\frac{2b}{\Delta},a_{12}=-b(1+\xi),a_{22}=\displaystyle 1-3b+2ab-\frac{b}{\Delta}.$
It can be seen that $b=|\alpha_{1}\alpha_{2}|=5,~~~~a=\displaystyle\frac{3b}{b+1}=\frac{5}{2},~~~~|1+\xi|^{2}=\displaystyle\frac{16}{5}$ and $\Delta=\left[(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}\right]=1.$
This gives, $a_{11}=14,~~a_{12}=-8-4i,~~~a_{22}=6.$
At this stage we applied Theorem 3.2 in the Case (2) above. However, here, though we have $D(\mu)=H(B)$, Theorem 3.2 is not applicable as it requires $\alpha_{1}$ and $\alpha_{2}$ to be distinct. As stated in [13, Remark 2.2], we need to suitably modify Theorem 3.2 in this case where $B$ has a pole of multiplicity $2$.
Here, we need to prove that Cauchy dual of $M_{z}$ on $D(\mu)$ is not subnormal. Therefore, we only need to obtain a condition that will allow us to conclude the same.
In view of [13, Theorem 3.5] it is enough to prove that for $l=1$, the $2\times 2$ principal minor of the matrix
(3.27) $\sum_{s=0}^{l}(-1)^{s}\binom{l}{s}((B_{m+1+s}B^{*}_{n+1+s}))_{m,n\geq 0}$
is negative.
A description of the above matrix and the computations leading to the negativity of the said minor have been carried out below.
We begin by stating a partial fraction formula which can be easily verified:
@@
Applying this to polynomials $p_{1}(z)/z$ and $p_{2}(z)/z$ we get,
@@
Note that $\frac{z}{z-\alpha}=-\displaystyle\sum_{l=0}^{\infty}\frac{z^{l+1}}{\alpha^{l+1}}\text{ and }\frac{z}{(z-\alpha)^{2}}=\sum_{l=0}^{\infty}\frac{(l+1)z^{l+1}}{\alpha^{l+2}}.$
@@
$b_{1}(z)$ $=\displaystyle\frac{p_{1}(z)}{q(z)}=\frac{p_{1}(z)}{(z-\alpha)^{2}}=\frac{p_{11}}{\alpha}\sum_{l=0}^{\infty}\frac{z^{l+1}}{\alpha^{l+1}}+\frac{p_{1}(\alpha)}{\alpha^{2}}\sum_{l=0}^{\infty}\frac{(l+1)z^{l+1}}{\alpha^{l+2}},$
(3.28) $b_{2}(z)$ $=\displaystyle\frac{p_{2}(z)}{q(z)}=\frac{p_{2}(\alpha)}{\alpha^{2}}\sum_{l=0}^{\infty}\frac{(l+1)z^{l+1}}{\alpha^{l+2}}.$
Let $b_{j}(z)=\displaystyle\sum_{m=1}^{\infty}b_{j,m}z^{m}$, $z\in\mathbb{D}$ and $j=1,2$. We note here that
$B_{m}=(b_{1,m},b_{2,m})$, $m\geq 1$.
Using equations (3.28) we get, $b_{1,m}=\displaystyle\frac{1}{\alpha^{m+1}}\left(p_{11}+\frac{p_{1}(\alpha)m}{\alpha^{2}}\right),~~m\geq 1$ and $b_{2,m}=\displaystyle\frac{p_{2}(\alpha)~m}{\alpha^{m+3}},~~m\geq 1.$ Thus, $\forall m\geq 0,s\geq 0$, we get,
$b_{1,m+1+s}=\displaystyle\frac{1}{\alpha^{m+2+s}}\left(p_{11}+\frac{p_{1}(\alpha)(m+1+s)}{\alpha^{2}}\right)$ and
$b_{2,m+1+s}=\displaystyle\frac{p_{2}(\alpha)~(m+1+s)}{\alpha^{m+4+s}}.$ After simplification we get,
@@
$N=\left|(n_{ij})_{0\leq i,j\leq 1}\right|$. Hence,
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{4}\overline{\alpha}^{2}}\left(1-\frac{2}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{4}\alpha^{2}}\left(1-\frac{2}{|\alpha|^{2}}\right),$
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{4}\overline{\alpha}^{3}}\left(2-\frac{3}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{6}\alpha}\left(1-\frac{2}{|\alpha|^{2}}\right),$
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{6}\overline{\alpha}}\left(1-\frac{2}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{4}\alpha^{3}}\left(2-\frac{3}{|\alpha|^{2}}\right),$
@@
$\hskip 14.22636pt+\frac{p_{11}\overline{p_{1}(\alpha)}}{|\alpha|^{6}\overline{\alpha}^{2}}\left(2-\frac{3}{|\alpha|^{2}}\right)+\frac{\overline{p_{11}}p_{1}(\alpha)}{|\alpha|^{6}\alpha^{2}}\left(2-\frac{3}{|\alpha|^{2}}\right).$
@@
$p_{11}=\sqrt{14},~p_{22}=\displaystyle\frac{2}{\sqrt{14}},~~p_{1}(\alpha)=\frac{20-30i}{\sqrt{14}}~~p_{2}(\alpha)=\frac{6+8i}{\sqrt{14}}.$
@@
$n_{00}=\displaystyle\frac{228}{625},~~~n_{01}=\frac{344+512i}{3125},~~~n_{10}=\frac{344-512i}{3125},~~~n_{11}=\frac{332}{3125}.$
Thus, $N=\displaystyle\frac{-2000}{9765625}<0.$ Hence, it follows that the Cauchy dual of $M_{z}$ on $D(\mu)$ is not subnormal.
Therefore, whenever measure $\mu=\delta_{\zeta_{1}}+\delta_{\zeta_{2}}$ where $\zeta_{1}$ and $\zeta_{2}$ are two distinct non-antipodal points on $\mathbb{T}$, the Cauchy dual of the operator $M_{z}$ on corresponding $D(\mu)$ is not subnormal. ∎
@@ 4. Epilogue
The authors believe that the main theorem of the present article can be proved for the measure $\mu$ in the general form, i.e. $\mu=c_{1}\delta_{\zeta_{1}}+c_{2}\delta_{\zeta_{2}},$ where $c_{1},c_{2}$ are positive real numbers, but with the techniques adopted here, the computational complexity increases substantially. In addition, it would be interesting to extend the main theorem when a measure $\mu$ is supported on three or more points on the unit circle. In particular, a suitable generalization of the notion of points on the unit circle being antipodal needs to be explored.
@@
Acknowledgments
The authors are thankful to Dr. Vaidehee Thatte and Dr. Mukul Sholapurkar for fruitful conversations during the preparation of this article.
Funding Declaration
This research has not received any funding.